I am trying to find the inverse equation of F(x)= (2-x)/(5x)
I get x=(2-y)/(5y)
multiply 5y through to get (5yx)=(2-y)
I divide the whole equation by y to get (5x)=(2/y)-(y/y) (y/y equaling 1)
so now my equation is 5x=(2/y)-1then take over the one to get 5x+1=2/y
I am not sure how to get y on its own after that please help thank you

Answer by Alan3354(69443)

I am trying to find the inverse equation of F(x)= (2-x)/(5x)
I get x=(2-y)/(5y)
==========================
5xy = 2-y
5xy + y = 2
y*(5x + 1) = 2
y = 2/(5x+1)